# Disentangling the Complex Multiplexed DIA Spectra in De Novo Peptide Sequencing

### v4

@@ 1 Introduction
Despite of their promising results, Traditional DDA methods suffer from limitations such as biased sampling and inconsistent detection of low-abundance peptides, leading to incomplete proteome coverage [11] [14] [15]. Data-Independent Acquisition (DIA) [15][16] addresses these limitations by fragmenting all ionized peptides within a predefined mass range, providing a more comprehensive and unbiased snapshot of the proteome [14].
DIA data fundamentally differ from DDA data in that DIA produces highly multiplexed spectra containing fragment ions from multiple coeluting peptides, where fragments of multiple peptides coexisting in the same spectrum [17] [18]. However, these characteristics come with significant challenges for de novo peptide sequencing. DIA spectra are inherently noisier due to the simultaneous fragmentation of all peptides, complicating the distinction between signal peaks and noise. The concurrent fragmentation leads to overlapping fragmentation patterns, making it difficult to assign fragment ions to specific peptides. The mixed signals from multiple coeluting peptides increase the complexity of spectral interpretation compared to DDA, where each spectrum typically associates with a single precursor ion. Furthermore, the need to process and analyze these complex, noisier spectra increases computational demands. In addition, DIA provides chromatogram information that offers a temporal profile of peptide ions, and the similarities between these chromatograms could help trace the source of the corresponding ion.
@@
We propose DIANovo, a framework specifically designed to address the complexities of DIA data introduced by coelution. Concretely, we first set to tackle the sheer size of highly-multiplexed DIA spectrum. Our encoder encodes the spectrum graph, like GraphNovo[13]. To reduce memory consumption, we implements with an automated edge generation process, which primarily focuses on learning the relative mass differences between graph nodes while the inter-node amino acid information is learned in a task-specific manner. To further reduce computational overhead, we leverage FlashAttention 2 [26], a linear-memory Transformer architecture designed for efficient computation. We also integrate dilated convolutional neural networks (CNNs)  [27] [28] to directly process chromatograms, enabling efficient encoding of time series information. Time-series embeddings are fed into Transformer, where the self-attention mechanism explicitly learns the similarities of the chromatograms, helping us to distinguish signal ions from those of coeluting peptides and noise. Moreover, we introduce a coelution-aware pretraining step that further improves model performance by incorporating coeluting peptides alongside the target peptide. This involves pretraining a model to predict ion types from coeluting peptides, with the resulting embeddings used as features in subsequent training stages. The coelution information helps the model better differentiate between signal and noise, leading to more accurate target peptide predictions. To the best of our knowledge, this is the first work to leverage coeluting peptide information in de novo sequencing, addressing a gap overlooked by previous methods.
@@ 2.1 De Novo Performance on Older-Generation Data
We compared our model’s performance with the baselines across several datasets [36] [37] [35] in Figure 1, where peptide recall is defined as the proportion of peptides whose entire sequences exactly match the database search results, without any mismatches or errors. On most datasets, our model demonstrated significantly better performance than the baselines, however the results are relatively lower across the board, indicating that there is a significant gap between database and de novo identification performance in older-generation mass spectrometers. Our amino acid recall is on average 60% higher than DeepNovo-DIA, while peptide recall being 53% higher. Comparing to Transformer DIA, our amino acid recall and peptide recall is 32% and 24% higher respectively.
@@ 2.3 Comparison of Peptide Detection by DDA and DIA
Figure 3: Comparison of Peptide Identification Under DDA or DIA Mode, with Orbitrap Q Exactive (older-generation).
@@
Figure 4: Comparison of Peptide Identification Under DDA or DIA Mode, with Orbitrap Astral.
@@ 2.5 Ablation Study
In this ablation study, illustrated in Figure 6, we evaluate the impact of different configurations on our model using the Yeast KO dataset. The results highlight that coelution-aware pretraining plays a helpful role in model performance, as its removal reduces peptide recall by 8.5% compared to DIANovo. More notably, the removal of Rotary Positional Encoding (RoPE) leads to a severe performance drop, with peptide recall plunging to 67.2% of the original. This decline suggests that absolute positional embeddings alone are insufficient for encoding graph node masses effectively. Additionally, excluding extended peak features or limiting the input to only three neighboring spectra results in moderate performance losses, demonstrating the importance of both extended features and broader spectral context. Finally, binning spectra into 0.5 Th intervals causes a significant drop in performance due to reduced spectral resolution, which increases ambiguity in peptide identification. We also included a version of our model without binning the spectrum. In this setting, we adopt a Casanovo[12]-style approach where raw peaks are encoded individually, and the retention time (RT) dimension is represented using sinusoidal positional embeddings. Our results show that although the no-bin version reduces preprocessing time, the binned model outperforms the no-bin variant significantly. We believe this improvement stems from the fact that the binned approach enables the model to better utilize the retention time dimension through its more structured and compact encoding. The degradation observed in these ablations underscores the essential role of these components in enhancing the model’s ability to identify peptides accurately.
@@ 4 Methods
Our model architecture is an encoder-decoder Transformer operating on the spectrum graph, including spectrum graph construction, dilated CNN (convolutional neural network) layers to process the chromatograms, RoPE (rotary positinal embedding)[40] integration to encode mass differences in the graph, coelution-aware pretraining to provide better understanding of the multiplexed spectra, and the two-stage decoding, where we predict the optimal path through the spectrum graph in stage 1, and refine the path to generate the final peptide sequence with mass tag filling in stage 2, in order to alleviate the sequence memorization problem commonly associated with de novo sequencing algorithms, where the model replies too much on previous seen peptide sequences during training. The model architecture is shown in Figure 9.
@@ 4.1 Feature Extraction
Finally, the spectrum graph is constructed by transforming every peak in the original spectrum from m/z to N-terminal residual masses. Each peak in the original spectrum is converted into 6 peaks in spectrum graph, considering 6 types of ions, including $a^{+}$, $a^{2+}$, $b^{+}$, $b^{2+}$, $y^{+}$, and $y^{2+}$, and we operate on the N-terminal residual masses, denoted as the graph node mass, following Equation 1, during which C-terminal ions are converted to their N-terminal residual masses by subtracting their C-terminal residual mass from the precursor mass. Note that although we did not specifically include other common ion types like $b-H_{2}O$ or $y-{NH_{3}}$, their peaks are converted into graph node as well, although they do not belong to the target sequence, their graph node can still contribute to our de novo objective by self attention. We provide an example of spectrum graph in Figure 10.
@@ 4.5 Coelution-aware Pretraining
A key difference between DDA and DIA data is that, in DDA, since precurosr selection is performed, there is typically a small number of peptides in the MS2 spectrum. Meanwhile in DIA, since we uniformly select a precursor range to perform fragmentation, there is a significant amount of coelution, i.e., one spectrum consisting of fragment ions of multiple coeluting peptide. In our study, coeluing peptides are identified by searching through the database search result, and include those whose RT (retention time) overlaps with the target de novo peptide, and their precursor m/z falls in the same isolation window as the target peptide.
@@ 4.6 Theoretical Analysis
    For real world test dataset, we substitute the experimental parameters with real data statistics, and takes into consideration the median peptide length, as well as median signal and noise intensities.
@@
    where #$\#\ De\ Novo\ Success$ and #$\#\ Database\ Success$ are the number of successfully identified successfully by de novo mode/ database search mode, $p$ is the p value, and #$\#\ De\ Novo\ Peptides$ and #$\#\ Database\ Peptides$ refer to the number of peptides with the same precursor mass to compare for de novo or database search, with #$\#\ De\ Novo\ Peptides\gg\#\ Database\ Peptides$
@@ Appendix D Precursor Feature Detection
For the detection of precursor features from liquid chromatography-mass spectrometry (LC-MS) maps, we utilized the same standard set of precursor information as employed by DeepNovo-DIA [19]. In our experiments, we implemented the detection results from DIA-NN [29]; however, these can be substituted with outputs from other existing peak detection algorithms, such as those referenced in Zhang et al. (2012) [38], Taynova et al. (2016) [45] or Tsou et al. (2015) [46]. The outcome of this detection step is a list of precursor features, each comprising the following essential information: feature ID, precursor mass-to-charge ratio (m/z), charge state, retention-time center, and scans across the retention-time range.
@@ Appendix G Model Implementation Details
For older-generation data, training and validation sets are Pain, PXD019777, and PXD003179, with 680,947/254,467/1,206,052 PSMs respectively. For Astral data, they are PXD046386 or PXD046453, with 1,896,662 and 1,086,577 PSMs respectively, these PSMs are chosen from the DIA-NN search result with 1% FDR.
We report the training and validation metrics for three datasets: the older-generation dataset, PXD046453, and PXD046386. The results are broken down by model component.
Older-Generation Dataset
  • Pretraining Model
    Metric Training Validation
    Cross-Entropy Loss 0.532 0.399
    Ion Accuracy (%) 82.8% 86.8%
  • Optimal Path Model
    Metric Training Validation
    Focal Loss 0.000134 0.000114
    Node Recall (%) 84.4% 87.3%
    Node Precision (%) 93.7% 94.7%
  • Sequence Generation Model
    Metric Training Validation
    Cross-Entropy Loss 0.337 0.335
PXD046453 Dataset
  • Pretraining Model
    Metric Training Validation
    Cross-Entropy Loss 0.422 0.312
    Ion Accuracy (%) 92.6% 92.6%
  • Optimal Path Model
    Metric Training Validation
    Focal Loss 0.000171 0.000124
    Node Recall (%) 74.8% 81.0%
    Node Precision (%) 89.9% 92.8%
  • Sequence Generation Model
    Metric Training Validation
    Cross-Entropy Loss 0.537 0.535
PXD046386 Dataset
  • Pretraining Model
    Metric Training Validation
    Cross-Entropy Loss 0.414 0.412
    Ion Accuracy (%) 87.2% 87.6%
  • Optimal Path Model
    Metric Training Validation
    Focal Loss 0.000100 0.0000944
    Node Recall (%) 87.4% 89.1%
    Node Precision (%) 93.9% 93.4%
  • Sequence Generation Model
    Metric Training Validation
    Cross-Entropy Loss 0.130 0.171
To prevent information leakage and ensure a rigorous evaluation of model generalization, we designed our training/validation/test splits to fully separate related peptide signals across sets, including coeluting peptides. Our model architecture leverages not only the target peptide but also contextual information from coeluting peptides within a spectrum. Therefore, to eliminate any direct or indirect sequence overlap, we adopted a raw file–based splitting strategy: for each training run, we designated one raw file exclusively for validation and used the remaining files for training. This guarantees that no peptides, or their coeluting partners, appear in both training and validation sets.
@@
For the test set, we ensured strict non-overlap in both peptide sequences and experimental conditions. All test sequences were drawn from datasets entirely distinct from those used in training. Furthermore, in our cross-species evaluation using Orbitrap Astral data, we imposed an even stricter criterion: training and testing data come from different species. Specifically, when training on the human dataset PXD046453, we tested only on the yeast dataset PXD046386. Conversely, when training on the yeast dataset PXD046386, we used entirely separate human datasets for testing. This species-based isolation helps ensure that learned representations do not exploit taxon-specific sequence features or instrument-specific noise patterns.
This design aligns with the DOME [50] guidelines by ensuring that spectra from the same peptide, protein, or coeluting group are never split across training and test partitions. All splits are reproducible and deterministic based on raw file boundaries.
@@ Appendix H Visualization of Learned Embeddings
We generate a t-SNE [51] plot from the learned peak embeddings, and observe that the model implicitly captures the relationship between coeluting precursors and their corresponding fragment ions, even when trained solely with the ion type loss. The t-SNE plot further reveals that peak embeddings are organized not only by fragment ion type (ion type label) but also by their source peptide (ion source label). For instance, in Supporting Figure 1, a dark blue cluster on the right side of the graph represents peaks originating from peptide 1 in the ion source plot. Within this cluster, the peaks are further subdivided into distinct colors (pink and green) in the ion type plot, corresponding to different fragment ion types. We provide the t-SNE plots from three different peptides to illustrate this effect.
@@
Appendix I Sensitivity Analysis
In this section, we present a sensitivity analysis on the raw file 20230320_OLEP08 _1000ngHeK_uPAC_180k-30min_MontBlanc _2p5ms_0p5_2Th_01 from the PXD046453 2p5ms dataset. Although this is a human raw file, we analyze it using the Mus musculus FASTA database for database search and perform de novo sequencing to assess how many peptides can be recovered beyond the database search results.
Using DIA-NN for database search, we identify 64,916 unique peptides from the Mus musculus FASTA. Performing de novo sequencing with feature detection, our algorithm produces 175,259 peptides, among which 20,408 peptides are correctly identified with full sequence accuracy, confirmed by searching the raw file against both the human and Mus musculus FASTA databases.
Within the Mus musculus database search results, 11,938 de novo-identified peptides match successfully, yielding an 18.4% peptide recall, consistent with previous observations. Additionally, outside the Mus musculus search results, we successfully identify 8,470 peptides exclusive to the human database.
This sensitivity analysis confirms that our de novo algorithm can accurately identify peptides beyond the scope of standard database search results.
List of supportingfigures 4 Venn diagram for the sensitivity analysis.
Appendix J Attention Pattern Analysis between RoPE and Absolute Positional Embeddings
To understand how positional encoding methods influence the modeling of fragment ions in de novo peptide sequencing, we compare rotary positional embeddings (RoPE) with absolute positional embeddings. Specifically, we examine how each method affects the model’s attention patterns and hidden representations of fragment ions.
Our analysis focuses on two aspects: attention scores between fragment ion pairs and cosine similarity between their hidden embeddings. Fragment ion pairs are categorized into four groups: ions from the same cleavage site, from different cleavage sites, one amino acid apart, and from distant amino acids.
As shown in Supporting Tables 5 and 6, RoPE consistently produces higher attention scores and embedding similarities across all categories. For ions from the same cleavage site, the attention score under RoPE is 0.166, compared to 0.119 with absolute PE. Similar trends hold for embedding similarity, where ions one amino acid apart reach 0.801 under RoPE, versus 0.523 with absolute PE. Even for distant ions, RoPE maintains stronger relationships, indicating a better ability to capture long-range dependencies.
These results suggest that RoPE encourages the model to focus more on functionally related ions and to represent them more coherently. This improved relational modeling likely enhances the model’s ability to interpret peptide fragmentation patterns effectively.
Category RoPE Attention Score Absolute PE Attention Score
Fragment ions from same cleavage site 0.119 0.166
Fragment ions from different cleavage sites 0.015 0.007
Fragment ions from one amino acid apart 0.054 0.010
Fragment ions from distant amino acids 0.002 0.001
List of supportingtables 5 Attention scores for different types of ions.
Category RoPE Embedding Similarity Absolute PE Emebdding Similarity
Fragment ions from same cleavage site 0.885 0.826
Fragment ions from different cleavage sites 0.719 0.454
Fragment ions from one amino acid apart 0.801 0.523
Fragment ions from distant amino acids 0.688 0.218
List of supportingtables 6 Embedding cosine similarities for different types of ions.
Appendix K Sensitivity to Coelution Number

### v5

@@ 1 Introduction
Despite their promising results, traditional DDA methods suffer from limitations such as biased sampling and inconsistent detection of low-abundance peptides, leading to incomplete proteome coverage [11] [14] [15]. Data-Independent Acquisition (DIA) [15][16] addresses these limitations by fragmenting all ionized peptides within a predefined mass range, providing a more comprehensive and unbiased snapshot of the proteome [14].
DIA data fundamentally differ from DDA data in that DIA produces highly multiplexed spectra containing fragment ions from multiple coeluting peptides, where fragments of multiple peptides coexist in the same spectrum [17] [18]. However, these characteristics come with significant challenges for de novo peptide sequencing. DIA spectra are inherently noisier due to the simultaneous fragmentation of all peptides, complicating the distinction between signal peaks and noise. The concurrent fragmentation leads to overlapping fragmentation patterns, making it difficult to assign fragment ions to specific peptides. The mixed signals from multiple coeluting peptides increase the complexity of spectral interpretation compared to DDA, where each spectrum typically associates with a single precursor ion. Furthermore, the need to process and analyze these complex, noisier spectra increases computational demands. In addition, DIA provides chromatogram information that offers a temporal profile of peptide ions, and the similarities between these chromatograms could help trace the source of the corresponding ion.
@@
We propose DIANovo, a framework specifically designed to address the complexities of DIA data introduced by coelution. Concretely, we first set to tackle the sheer size of highly-multiplexed DIA spectrum. Our encoder encodes the spectrum graph, like GraphNovo[13]. To reduce memory consumption, we implements an automated edge generation process, which primarily focuses on learning the relative mass differences between graph nodes while the inter-node amino acid information is learned in a task-specific manner. To further reduce computational overhead, we leverage FlashAttention 2 [26], a linear-memory Transformer architecture designed for efficient computation. We also integrate dilated convolutional neural networks (CNNs)  [27] [28] to directly process chromatograms, enabling efficient encoding of time series information. Time-series embeddings are fed into the Transformer encoder, where the self-attention mechanism explicitly learns the similarities of the chromatograms, helping us to distinguish signal ions from those of coeluting peptides and noise. Moreover, we introduce a coelution-aware pretraining step that further improves model performance by incorporating coeluting peptides alongside the target peptide. This involves pretraining a model to predict ion types from coeluting peptides, with the resulting embeddings used as features in subsequent training stages. The coelution information helps the model better differentiate between signal and noise, leading to more accurate target peptide predictions. To the best of our knowledge, this is the first work to leverage coeluting peptide information in de novo sequencing, addressing a gap overlooked by previous methods.
@@ 2.1 De Novo Performance on Older-Generation Data
We compared our model’s performance with the baselines across several datasets [36] [37] [35] in Figure 1, where peptide recall is defined as the proportion of peptides whose entire sequences exactly match the database search results, without any mismatches or errors. On most datasets, our model demonstrated significantly better performance than the baselines, however the results are relatively lower across the board compared to DDA levels [2] [13], indicating that there is a significant gap between database and de novo identification performance in older-generation mass spectrometers. Our amino acid recall is on average 60% higher than DeepNovo-DIA, while peptide recall being 53% higher. Comparing to Transformer DIA, our amino acid recall and peptide recall is 32% and 24% higher respectively.
@@ 2.3 Comparison of Peptide Detection by DDA and DIA
Figure 3: Venn diagram, comparison of peptide identification under DDA or DIA mode, with Orbitrap Q Exactive (older-generation), where blue and orange circles refer to number of peptides identified in database search, while pink and green circles refer to number of peptides identified in de novo mode, under DDA and DIA respectively.
@@
Figure 4: Venn diagram, comparison of peptide identification Under DDA or DIA mode, with Orbitrap Astral.
@@ 2.5 Ablation Study
In this ablation study, illustrated in Figure 6, we evaluate the impact of different configurations on our model using the Yeast KO dataset. The results highlight that coelution-aware pretraining plays a helpful role in model performance, as its removal reduces peptide recall by 8.5% compared to DIANovo. More notably, the removal of Rotary Positional Encoding (RoPE) leads to a severe performance drop, with peptide recall plunging to 67.2% of the original. This decline suggests that absolute positional embeddings alone are insufficient for encoding graph node masses effectively. Additionally, excluding extended peak features or limiting the input to only three neighboring spectra results in moderate performance losses, demonstrating the importance of both extended features and broader spectral context. Finally, binning spectra into 0.5 Th intervals causes a significant drop in performance due to reduced spectral resolution, which increases ambiguity in peptide identification.
@@ 4 Methods
Our model architecture is an encoder-decoder Transformer operating on the spectrum graph, including spectrum graph construction, dilated CNN (convolutional neural network) layers to process the chromatograms, RoPE (rotary positinal embedding)[40] integration to encode mass differences in the graph, coelution-aware pretraining to provide better understanding of the multiplexed spectra, and the two-stage decoding, where we predict the optimal path through the spectrum graph in stage 1, and refine the path to generate the final peptide sequence with mass tag filling in stage 2, in order to alleviate the sequence memorization problem commonly associated with de novo sequencing algorithms, where the model relies too much on previous seen peptide sequences during training. The model architecture is shown in Figure 9.
@@ 4.1 Feature Extraction
Finally, the spectrum graph is constructed by transforming every peak in the original spectrum from m/z to N-terminal residual masses. Each peak in the original spectrum is converted into 6 peaks in spectrum graph, taking into account 6 types of ions, including $a^{+}$, $a^{2+}$, $b^{+}$, $b^{2+}$, $y^{+}$, and $y^{2+}$. We operate on the N-terminal residual masses, denoted as the graph node mass, following Equation 1, during which C-terminal ions are converted to their N-terminal residual masses by subtracting their C-terminal residual mass from the precursor mass. Note that although we did not specifically include other common ion types like $b-H_{2}O$ or $y-{NH_{3}}$, their peaks are converted into graph node as well. They do not belong to the target sequence, but their graph node can still contribute to our de novo objective by self attention. We provide an example of spectrum graph in Figure 10.
@@ 4.5 Coelution-aware Pretraining
A key difference between DDA and DIA data is that, in DDA, since precurosr selection is performed, there is typically a small number of peptides in the MS2 spectrum. Meanwhile in DIA, since we uniformly select a precursor range to perform fragmentation, there is a significant amount of coelution, i.e., one spectrum consisting of fragment ions of multiple coeluting peptides. In our study, coeluing peptides are identified by searching through the database search result, and include those whose RT (retention time) overlaps with the target de novo peptide, and their precursor m/z falls in the same isolation window as the target peptide.
@@ 4.6 Theoretical Analysis
    For real world test dataset, we substitute the experimental parameters with real data statistics, and take into consideration the median peptide length, as well as median signal and noise intensities.
@@
    where #$\#\ De\ Novo\ Success$ and #$\#\ Database\ Success$ are the number of successfully identified peptides by de novo mode/ database search mode, $p$ is the p value, and #$\#\ De\ Novo\ Peptides$ and #$\#\ Database\ Peptides$ refer to the number of peptides with the same precursor mass to compare for de novo or database search, with #$\#\ De\ Novo\ Peptides\gg\#\ Database\ Peptides$
@@ Appendix D Precursor Feature Detection
For the detection of precursor features from liquid chromatography-mass spectrometry (LC-MS) maps, we utilized the same standard set of precursor information as employed by DeepNovo-DIA [19]. In our experiments, we implemented the detection results from DIA-NN [29] during the main experiment, and DIAUmpire [45] during the comparison with Cascadia; however, these can be substituted with outputs from other existing peak detection algorithms, such as those referenced in Zhang et al. (2012) [38], Taynova et al. (2016) [46] or Tsou et al. (2015) [45]. The outcome of this detection step is a list of precursor features, each comprising the following essential information: feature ID, precursor mass-to-charge ratio (m/z), charge state, retention-time center, and scans across the retention-time range.
@@ Appendix G Model Implementation Details
For older-generation data, training and validation sets are Pain, PXD019777, and PXD003179, with 680,947/254,467/1,206,052 PSMs respectively. For Astral data, they are PXD046386 or PXD046453 + PXD046453, with 1,896,662, 1,086,577, and 2,136,215 PSMs respectively, these PSMs are chosen from the DIA-NN search result with 1% FDR.
To prevent information leakage and ensure a rigorous evaluation of model generalization, we designed our training/validation/test splits to fully separate related peptide signals across sets, including coeluting peptides. Our model architecture leverages not only the target peptide but also contextual information from coeluting peptides within a spectrum. Therefore, to eliminate any direct or indirect sequence overlap, we adopted a raw file–based splitting strategy: for each training run, we designated a specific RT range in one raw file exclusively for validation and used the remaining for training. This guarantees that no peptides, or their coeluting partners, appear in both training and validation sets.
@@
For the test set, we ensured strict non-overlap in both peptide sequences and experimental conditions. All test sequences were drawn from datasets entirely distinct from those used in training. Furthermore, in our cross-species evaluation using Orbitrap Astral data, we imposed an even stricter criterion: training and testing data come from different species. Specifically, when training on the human datasets PXD046453 + PXD046444, we tested only on the yeast dataset PXD046386. Conversely, when training on the yeast dataset PXD046386, we used entirely separate human datasets for testing.
@@ Appendix H Visualization of Learned Embeddings
We generate a t-SNE [50] plot from the learned peak embeddings, and observe that the model implicitly captures the relationship between coeluting precursors and their corresponding fragment ions, even when trained solely with the ion type loss. The t-SNE plots (shown in Supporting Figure 1, 2, and 3) further reveal that peak embeddings are organized not only by fragment ion type (ion type label) but also by their source peptide (ion source label). For instance, in Supporting Figure 1, a dark blue cluster on the right side of the graph represents peaks originating from peptide 1 in the ion source plot. Within this cluster, the peaks are further subdivided into distinct colors (pink and green) in the ion type plot, corresponding to different fragment ion types. We provide the t-SNE plots from three different peptides to illustrate this effect.
@@
Appendix I Sensitivity to Coelution Number